OpenStudy (anonymous):

what is the length of the base on a square pyramid ? if the H=3 cm and V=36 cm

OpenStudy (anonymous):

hold on is it 12 right

ganeshie8 (ganeshie8):

how did u get 12 ?

OpenStudy (anonymous):

36/3

ganeshie8 (ganeshie8):

thats a good try, but no ! here is the formula for volume of pyramid :\[\large V = \dfrac{1}{3}Bh \]

ganeshie8 (ganeshie8):

plugin the given numbers and solve \(B\) first

ganeshie8 (ganeshie8):

\[\large 36= \dfrac{1}{3}B(3) \]

ganeshie8 (ganeshie8):

\[\large 36= B\]

ganeshie8 (ganeshie8):

since it is a square pyramid, the base will be a square whats the formula for area of square ?

OpenStudy (anonymous):

width times height

OpenStudy (anonymous):

Formula: Width × Height

ganeshie8 (ganeshie8):

\[\large 36 = s^2\] solve \(\large s\)

OpenStudy (anonymous):

so it would be 8 cm

OpenStudy (anonymous):

hold on

OpenStudy (anonymous):

still not getting it ahhhh

OpenStudy (anonymous):

6 cm

ganeshie8 (ganeshie8):

thats right ! length of side of square base = 6cm

OpenStudy (anonymous):

yes thank you so much I am so frustrated have final exam in morning and its already 3 am here

ganeshie8 (ganeshie8):

you will do great, don't wry :) you need to sleep NOW !